Calculate the pH

pH of a solution in which one normal adult dose of aspirin (650 mg

mg ) is dissolved in 6.0 ounces of water.


1
Expert's answer
2022-02-28T11:03:50-0500

Molar mass of C9H8O4 is 180.16 g/mol 

9 ounce [US, liquid] = 266mL 

640mg = 0.640g 

I work on the basis that the final volume is 9 ounces = 266mL 

Mol C9H8O4 in 0.640g = 0.640/180.16 = 0.00355 mol 

Molarity of solution = 0.00355/0.266 = 0.0133M 

If pKa = 3.5 , Ka = 10^-3.5 = 3.16*10^-4 

Calculate [H+] 

Ka = [H+]² / [acid] 

3.16*10^-4 = [H+]² / 0.0133 

[H+]² = (3.16*10^-4)* 0.0133 

[H+]² = 4.20*10^- 6 

[H+] = 2.05*10^-3M 

pH = -log (2.05*10^-3) 

pH = 2.69